# An NV- center in magnesium oxide as a spin qubit for hybrid quantum technologies

### v1

@@ Abstract
Recent predictions suggest that oxides, such as MgO and CaO, could serve as hosts of spin defects with long coherence times and thus be promising materials for quantum applications. However, in most cases specific defects have not yet been identified. Here, by using a high-throughput first-principles framework and advanced electronic structure methods, we identify an NV--like center in MgO with favorable electronic and optical properties for hybrid quantum technologies. We show that this defect has stable triplet ground and excited states, with singlet shelving states enabling optical initialization and spin-dependent readout. We predict several properties, including absorption, emission, and zero-phonon line energies, as well as zero-field splitting tensor, and hyperfine interaction parameters, which can aid in the experimental identification of this defect. Our calculations show that due to a strong pseudo-Jahn Teller effect and low frequency phonon modes, the NV- center in MgO is subject to a substantial vibronic coupling. We discuss design strategies to reduce such coupling and increase the Debye-Waller factor, including the effect of strain and the localization of the defect states. We propose that the favorable properties of the NV- defect, along with the technological maturity of MgO, could enable hybrid classical-quantum applications, such as spintronic quantum sensors and single qubit gates.
@@ I Introduction
The nitrogen-vacancy center (NV-) in diamond [1] and the divacancy (VV0) in silicon carbide (SiC) [2] are prototypical spin defects for quantum technologies, including sensing, communication, and computing. A key requirement for these applications is the long coherence time of the spin defect as well as scalability and integration of the host into current microelectronic platforms. However, the coherence time of defects in diamond and SiC is intrinsically limited by the presence of non-zero nuclear spin isotopes (namely 13C and 29Si, respectively). Additionally, achieving high-yield, scalable fabrication of diamond devices is challenging due to difficulties in etching and polishing diamond [3], and the synthesis of high-quality thin films of a single phase of SiC is a challenge due to its many nearly-degenerate polytypes [4]. Therefore, the identification of new host materials for spin defects is an active field of research.
@@
In this work, we identify and characterize a promising spin defect in MgO, which has a long and rich history in the microelectronics industry, particularly in spintronics. Magnesium oxide plays a key role in enabling giant tunnel magnetoresistance in magnetic tunnel junctions used in random access memories [7, 8, 9, 10, 11]. In addition, MgO tunnel barriers enhance spin injection efficiency from the ferromagentic source to the semiconducting channel in spin field effect transistors [12]. Although controversial, it has been suggested that MgO might exhibit d0-ferromagnetism [13, 14, 15], which could enable the use of the material as a dilute magnetic semiconductor spin filter in spin transistors. We note that the high dielectric constant of MgO, its low leakage current, and thermal stability make it an ideal material for use as a gate dielectric in thin film transistors [16]. Furthermore, MgO is CMOS-compatible and can be grown in the form of high quality thin films via a variety of methods [17, 18], and it is frequently used as a template layer for ferroelectrics [19] and superconductors [20]. Thus, identifying spin defects in MgO could open avenues to combine spintronics, ferroelectrics, and quantum information science to realize multifunctional, hybrid classical-quantum applications.
@@ II.1 High-throughput screening of spin defects in MgO
In our search, we considered isolated defects and defect complexes (the latter comprising two neighboring individual defects), including native vacancies and interstitials and their complexes, and substitutional and interstitial defects and defect complexes with $s$- and $p$-elements of the periodic table. We limited our investigation to $s$- and $p$-dopants as we use DFT at the PBE level to screen defects, and we expect that $d$- and $f$-dopants would require the use of DFT+U or hybrid DFT methods, which will be the topic of future investigations, given their higher computational cost. For simplicity, we excluded interstitial-interstitial complexes from the screening. We also restricted our search to complexes with first or second nearest-neighbor defect constituents (i.e. the maximum distance between the two constituent defects in a complex is 3.5 $\mathrm{\AA}$), given the cell sizes that are reasonably affordable in our high-throughput procedure, and we included a single type of extrinsic dopant (i.e. we do not consider defect complexes of the type $\mathrm{X_{Mg}Y_{Mg}}$, $\mathrm{X_{Mg}Y_{O}}$, $\mathrm{X_{Mg}Int_{Y}}$, $\mathrm{X_{O}Int_{Y}}$, where X and Y are different elements substituting the Mg or O site or existing as an interstitial (Int)). Our initial screening yielded 2917 defects, of which 1088 were determined to be on the defect hull–i.e. to have the lowest formation energy per stoichiometry and per Fermi energy [25]. Out of those, we considered the defects with a stable spin triplet ground state (which can thus serve as a two-level system) and which possess a zero-phonon line (ZPL), arriving at the list of the 40 defects tabulated in the Supplementary Information (SI).
@@
Figure 1: High-throughput screening of spin defects in MgO. Schematic of the workflow, which includes creation of isolated defects and complexes including $s$- and $p$-dopants and intrinsic defects, which are screened based on stability, presence of a spin triplet ground state and zero-phonon line, and synthesizability (see text), leading to the identification of an NV--like center in MgO.
@@
We found that 24 of the 40 defects are complexes comprised of second nearest-neighbor defects, which are excluded from further analysis as their experimental synthesis/positioning will likely be challenging. The remaining isolated defects and first nearest-neighbor complexes are shown in Table 1. We find that $\mathrm{Ga_{O}}$ is the only isolated defect with a stable spin triplet ground state; however, we did not consider it since $\mathrm{Ga_{Mg}}$ has a much lower formation energy and thus it is more likely to be formed when $\mathrm{Ga}$ is implanted in MgO. Several complexes listed in Table 1 are of the type $\mathrm{X_{O}Vac_{Mg}}$, where the dopant X (X = B, I, Sb, Bi) lies between the magnesium and oxygen vacancy sites. This class of defects was first identified in CaO, where the Bi dopant has been proposed to be suitable for quantum applications due to the presence of a clock transition [6]. These defects were excluded from the present study, as we expect them to have similar optical and spin properties as those in CaO, including, unfortunately, a low Debye-Waller factor. Note that in MgO, the $\mathrm{B_{O}Vac_{Mg}}$ defect is stable, unlike in CaO, but it has a higher formation energy (by  7 eV) than those discussed below. Further, several complexes of the type $\mathrm{X_{Mg}X_{O}}$, where X = B, Al, Ga, In, Tl are stable, with a detectable ZPL; however they are likely difficult to synthesize due to the slow diffusion of substitutional impurities [24] and high formation energies (due to cations substituting the oxygen site), and hence they were not furthered analyzed. Interestingly, our search identified additional five defects with a spin triplet ground state and possessing a ZPL; but they all consist of anions substituting the magnesium site and/or cations substituting the oxygen site, and thus they will likely have large formation energies and be challenging to synthesize (see SI).
@@
Finally, we are left with the defect $\mathrm{Int_{N}Vac_{Mg}}$, comprising a nitrogen interstitial next to a magnesium vacancy; it is similar to the NV- in diamond, which however is formed by a nitrogen substitutional, instead of interstitial, next to a carbon vacancy. The synthesis of $\mathrm{Int_{N}Vac_{Mg}}$ is expected to be straight-forward, since interstitials generally have low diffusion barriers in MgO [26, 27]. Moreover, nitrogen doping of MgO via, e.g. molecular beam epitaxy and implantation has been reported in the literature [28, 29, 30, 31, 32, 33]. Thus, we choose to investigate the $\mathrm{Int_{N}Vac_{Mg}}$ defect (referred to as NV- from here onwards) for the remainder of our study.
@@ II.2 Ground state properties using hybrid DFT
We investigated the ground state properties of the NV- center in MgO using density functional theory (DFT) with the dielectric-dependent hybrid (DDH) exchange-correlation functional [41]. The inverse of the experimental macroscopic dielectric constant $\epsilon_{\infty}=2.96$ was used as the fraction of exact exchange, i.e. $\alpha$ = 0.34 [42, 41]. The DDH functional is able to recover the experimental band gap at 0 K accurately (after accounting for zero-point renormalization effects), unlike semi-local functionals like PBE, which underestimate the bandgap (see SI Table I for a comparison between the band gaps obtained using PBE, DDH, and experiment).
The NV- defect in MgO has $C_{4v}$ symmetry in the ground state, with a nitrogen atom present at the bond-center interstitial site, bonded to four neighboring magnesium atoms (dN-Mg = 2.49 Å) and an oxygen atom (dN-O = 1.33 Å), with a nearest neighbor magnesium vacancy and an additional electron. The fourfold rotation axis ($C_{4}$) is along the N-O bond in the $<$001$>$ direction. We plot the defect formation energies of relevant defects as a function of the Fermi level in Fig. 2a. We see that the NV- is stable within the bandgap, and therefore, may be experimentally realizable. We expect that NV- will be compensated by positively charged oxygen vacancies.
@@
Overall, the ground state properties of the NV- in MgO indicate that this defect is stable over a large energy range within the band gap of the host and thus, should be experimentally realizable. It has a spin triplet ground state with isolated occupied and unoccupied levels, pointing to the possibility of optical transitions to prepare and measure a qubit state. The computed ZFS tensor shows that a sufficiently large separation between spin sublevels is present, and that initialization in the $m_{s}$ = 0 sublevel should be realizable experimentally. Further, the computed hyperfine parameters indicate that the NV- center should be distinguishable from other nitrogen defects in MgO.
@@ II.3 Excited state properties using hybrid TDDFT
We now turn to investigate the excited state properties of the NV- center using time-dependent density functional theory (TDDFT) with the DDH functional, as implemented in the WEST code [47]. We used spin-conserving and spin-flip TDDFT-DDH to investigate triplet and singlet excited states, respectively. The relaxed structure of the triplet ground state and first excited triplet state (obtained using DDH and TDDFT-DDH respectively) are shown in Fig. 3a.
@@
Having determined the vertical excitation (absorption) energies, we proceeded to relax the structure in the first excited triplet (${}^{3}E$) state using TDDFT-DDH with analytical forces [47] (see Fig. 3a). A decomposition of the many-body state shows that the $a_{1}$rightarrow $e$ excitations are dominant. We note that in the relaxed geometry of the excited state, the symmetry is reduced to $C_{1}$, with dN-Mg = 1.98 - 2.86 Åand dN-O = 1.37 Å. We find four equivalent minima, with the N-O bond tilted along the $+x$, $-x$, $+y$, $-y$ directions, consistent with the $C_{4v}$ symmetry of the ground state. Table 3 lists the absorption, zero phonon line (ZPL), and emission energies of the NV- in MgO (calculated under the Franck-Condon approximation) and compares it with that of the NV- center in diamond and the VV0 in silicon carbide. In MgO, the absorption and ZPL of the NV- center are in the ultra-violet (UV) range whereas the emission is in the visible range, highlighting a large Stokes shift.
@@
A large Stokes shift implies that a large number of phonons are involved in the emission process, which could lower the intensity of the ZPL relative to the rest of the emission spectrum, thus yielding an unfavourable Debye-Waller factor (DWF). The average number of phonons emitted during an electronic transition is called the Huang-Rhys factor (HRF, denoted as $S$) and is given by [54]
@@
We calculated $\omega_{eff}$, and therefore, the HRF $S$, by constructing a one-dimensional configuration coordinate diagram (CCD) along the direction connecting the relaxed ground and excited state structures [56, 53] (geometry optimized using DDH and TDDFT-DDH, respectively), as shown in Fig. 3c. By fitting a quadratic function around the points near the respective minima, we obtain the effective phonon frequencies of the ground and excited states to be $\hbar\omega_{g}$ = 32.79 meV and $\hbar\omega_{e}$ = 61.77 meV respectively. Using $\hbar\omega_{g}$ and $\Delta Q$, we obtain the HRF for emission as $S$ = 44.49, which is again indicative of a strong electron-phonon coupling [54], as expected. Such a large HRF leads to a negligible Debye-Waller factor (DWF $\simeq e^{-S}$approx 10-20), as a majority of the photons contribute to the emission side band instead of the ZPL. Thus, the ZPL will likely be undetectable, rendering the defect unsuitable as a single photon emitter (for example, for quantum communication applications). However, the defect’s triplet ground state, large ZFS, and many-body level structure may make it suitable for quantum sensing, computing, or transduction, by enabling optical initialization and spin-selective photo-luminescence. Moreover, nanostructuring or the additional presence of extended defects can be used to lower the HRF and therefore increase the intensity of the ZPL, as explored in SiC [57].
Using $\hbar\omega_{g}$, $\hbar\omega_{e}$, and $\Delta Q$, we show the emission spectrum of the NV- in Fig. 3d, which is nearly a Gaussian with a peak at the Franck-Condon energy of 1.43 eV, and a line-width of 420 meV. The emission spectrum thus lies in the visible red-infrared region, making it easy to detect. Moreover, a broad line-width is advantageous as it would allow multiple frequency filters to be used (for example, to filter out emissions from an underlying substrate) while retaining a significant fraction of the emission spectrum to be used for detection.
@@
To summarize the computed excited state properties of the NV- in MgO, we find the presence of triplet excited and singlet shelving states, pointing at the possibility of an optical initialization cycle and spin-selective photo-luminescence. Further, our results show that the absorption and ZPL energies are in the UV range, whereas the emission energy is in the visible (red) range. The visible emission energy, together with the computed broad emission side-band, is a convenient feature for experimental detection in applications such as quantum sensing and transduction. We also find a large $\Delta Q$ and related Stokes shift, large HRF, and low DWF, indicating strong vibronic coupling. Identifying the reason behind this strong coupling is key to determining engineering strategies that can mitigate it and to enhance our understanding of oxides as hosts for spin defects. We explore this in further detail below.

### v2

@@ Abstract
Recent predictions suggest that oxides, such as MgO and CaO, could serve as hosts of spin defects with long coherence times and thus be promising materials for quantum applications. However, in most cases specific defects have not yet been identified. Here, by using a high-throughput first-principles framework and advanced electronic structure methods, we identify a negatively-charged complex between a nitrogen interstitial and a magnesium vacancy in MgO with favorable electronic and optical properties for hybrid quantum technologies. We show that this NV- center has stable triplet ground and excited states, with singlet shelving states enabling optical initialization and spin-dependent readout. We predict several properties, including absorption, emission, and zero-phonon line energies, as well as zero-field splitting tensor, and hyperfine interaction parameters, which can aid in the experimental identification of this defect. Our calculations show that due to a strong pseudo-Jahn Teller effect and low frequency phonon modes, the NV- center in MgO is subject to a substantial vibronic coupling. We discuss design strategies to reduce such coupling and increase the Debye-Waller factor, including the effect of strain and the localization of the defect states. We propose that the favorable properties of the NV- defect, along with the technological maturity of MgO, could enable hybrid classical-quantum applications, such as spintronic quantum sensors and single qubit gates.
@@ I Introduction
The nitrogen-vacancy center (NV-) in diamond [1] and the divacancy (VV0) in silicon carbide (SiC) [2] are prototypical spin defects for quantum technologies, including sensing, communication, and computing. A key requirement for these applications is the long coherence time of the spin defect as well as the scalability and integration of the host into current microelectronic platforms. However, the coherence time of defects in diamond and SiC is intrinsically limited by the presence of non-zero nuclear spin isotopes (namely 13C and 29Si, respectively). Additionally, achieving high-yield, scalable fabrication of diamond devices is challenging due to difficulties in etching and polishing diamond [3], and the synthesis of high-quality thin films of a single phase of SiC is a challenge due to its many nearly-degenerate polytypes [4]. Therefore, the identification of new host materials for spin defects is an active field of research.
@@
In this work, we identify and characterize a promising spin defect in MgO, which has a long and rich history in the microelectronics industry, particularly in spintronics. Magnesium oxide plays a key role in enabling giant tunnel magnetoresistance in magnetic tunnel junctions used in random access memories [7, 8, 9, 10, 11]. In addition, MgO tunnel barriers enhance spin injection efficiency from the ferromagentic source to the semiconducting channel in spin field effect transistors [12]. Although controversial, it has been suggested that MgO might exhibit d0-ferromagnetism [13, 14, 15], which could enable the use of the material as a dilute magnetic semiconductor spin filter in spin transistors. We note that the high dielectric constant of MgO, its low leakage current, and thermal stability make it an ideal material for use as a gate dielectric in thin film transistors [16]. Furthermore, MgO is CMOS-compatible and can be grown in the form of high quality thin films via a variety of methods [17, 18], and it is frequently used as a template layer for ferroelectrics [19, 20] and superconductors [21]. Thus, identifying spin defects in MgO could open avenues to combine spintronics, ferroelectrics, and quantum information science for the realization of multifunctional, hybrid classical-quantum applications.
@@ II.1 High-throughput screening of spin defects in MgO
In our search, we considered isolated defects and defect complexes (the latter comprising two neighboring individual defects), including native vacancies and interstitials and their complexes, and substitutional and interstitial defects and defect complexes with $s$- and $p$-elements of the periodic table. We limited our investigation to $s$- and $p$-dopants as we use DFT at the PBE level to screen defects, and we expect that $d$- and $f$-dopants would require the use of DFT+U or hybrid DFT methods, which will be the topic of future investigations, given their higher computational cost. For simplicity, we excluded interstitial-interstitial complexes from the screening. We also restricted our search to complexes with first or second nearest-neighbor defect constituents (i.e. the maximum distance between the two constituent defects in a complex is 3.5 $\mathrm{\AA}$), given the cell sizes that are reasonably affordable in our high-throughput procedure, and we included a single type of extrinsic dopant (i.e. we do not consider defect complexes of the type $\mathrm{X_{Mg}Y_{Mg}}$, $\mathrm{X_{Mg}Y_{O}}$, $\mathrm{X_{Mg}Int_{Y}}$, $\mathrm{X_{O}Int_{Y}}$, where X and Y are different elements substituting the Mg or O site or existing as an interstitial (Int)). Our initial screening yielded 2917 defects, of which 1088 were determined to be on the defect hull–i.e. to have the lowest formation energy per stoichiometry and per Fermi energy [26]. Out of those, we considered the defects with a stable spin triplet ground state (which can thus serve as a two-level system) and which possess a zero-phonon line (ZPL), arriving at the list of the 40 defects tabulated in Supplementary Tables 1 and 2.
@@
Figure 1: High-throughput screening of spin defects in MgO. Schematic of the workflow, which includes creation of isolated defects and complexes including $s$- and $p$-dopants and intrinsic defects, which are screened based on stability, presence of a spin triplet ground state and zero-phonon line, and synthesizability (see text), leading to the identification of an NV- center in MgO (shown in inset).
@@
We found that 24 of the 40 defects are complexes comprised of second nearest-neighbor defects, which are excluded from further analysis as their experimental synthesis/positioning will likely be challenging. The remaining isolated defects and first nearest-neighbor complexes are shown in Table 1. We find that $\mathrm{Ga_{O}}$ is the only isolated defect with a stable spin triplet ground state; however, we did not consider it since $\mathrm{Ga_{Mg}}$ has a much lower formation energy and thus it is more likely to be formed when $\mathrm{Ga}$ is implanted in MgO. Several complexes listed in Table 1 are of the type $\mathrm{X_{O}Vac_{Mg}}$, where the dopant X (X = B, I, Sb, Bi) lies between the magnesium and oxygen vacancy sites. This class of defects was first identified in CaO, where the Bi dopant has been proposed to be suitable for quantum applications due to the presence of a clock transition [6]. These defects were excluded from the present study, as we expect them to have similar optical and spin properties as those in CaO, including, unfortunately, a low Debye-Waller factor. Note that in MgO, the $\mathrm{B_{O}Vac_{Mg}}$ defect is stable, unlike in CaO, but it has a higher formation energy (by  7 eV) than those discussed below. Further, several complexes of the type $\mathrm{X_{Mg}X_{O}}$, where X = B, Al, Ga, In, Tl are stable, with a detectable ZPL; however they are likely difficult to synthesize due to the slow diffusion of substitutional impurities [25] and high formation energies (due to cations substituting the oxygen site), and hence they were not further analyzed. Interestingly, our search identified additional five defects with a spin triplet ground state and possessing a ZPL; but they all consist of anions substituting the magnesium site and/or cations substituting the oxygen site, and thus they will likely have large formation energies and be challenging to synthesize (see Supplementary Note 1).
@@
Finally, we are left with the defect $\mathrm{Int_{N}Vac_{Mg}}$, comprising a nitrogen interstitial next to a magnesium vacancy; it is similar to the NV- in diamond, which however is formed by a nitrogen substitutional, instead of interstitial, next to a carbon vacancy. The synthesis of $\mathrm{Int_{N}Vac_{Mg}}$ is expected to be straight-forward, since interstitials generally have low diffusion barriers in MgO [27, 28]. Moreover, nitrogen doping of MgO via, e.g. molecular beam epitaxy and implantation has been reported in the literature [29, 30, 31, 32, 33, 34]. Thus, we choose to investigate the $\mathrm{Int_{N}Vac_{Mg}}$ defect (referred to as NV- hereafter) for the remainder of our study.
@@ II.2 Ground state properties using hybrid DFT
We investigated the ground state properties of the NV- center in MgO using density functional theory (DFT) with the dielectric-dependent hybrid (DDH) exchange-correlation functional [42]. The inverse of the experimental macroscopic dielectric constant $\epsilon_{\infty}=2.96$ was used as the fraction of exact exchange, i.e. $\alpha$ = 0.34 [43, 42]. The DDH functional is able to recover the experimental band gap at 0 K accurately (after accounting for zero-point motion renormalization effects), unlike semi-local functionals like PBE, which underestimate the bandgap (see Supplementary Table 3 for a comparison between the band gaps obtained using PBE, DDH, and experiment).
As shown in the inset of Fig. 1, the NV- defect in MgO has $C_{4v}$ symmetry in the ground state, with a nitrogen atom present at the bond-center interstitial site, bonded to four neighboring magnesium atoms (dN-Mg = 2.49 Å) and an oxygen atom (dN-O = 1.33 Å), with a nearest neighbor magnesium vacancy and an additional electron. The fourfold rotation axis ($C_{4}$) is along the N-O bond in the $<$001$>$ direction. We plot the defect formation energies of relevant defects under O2-rich, N2-rich conditions as a function of the Fermi level in Fig. 2a. We see that the NV- is stable within the bandgap, and therefore, may be experimentally realizable. Further, we expect that NV- will be compensated by positively charged oxygen vacancies. Note that the Fermi energy may be pinned at 2.3 eV above the valence band maximum (VBM) due to compensating native defects (magnesium and oxygen vacancies) under high temperature equilibrium growth conditions. One strategy to move the equilibrium Fermi level into the stability region of the NV- defect could be to $n$-dope MgO during growth, for example, with Al3+ [44] or Fe3+ [45]. This would have to be done carefully to avoid the introduction of additional optically active defects and a high concentration of spinful nuclei into the sample. However, sample preparation techniques could be designed to circumvent the requirement of $n$-type doping. Specifically, sample preparation typically involves high temperature growth in a nitrogen-rich environment, followed by a cooling down process where the nitrogen flux is turned off, thereby fixing the concentration of nitrogen in the sample. Under these low temperature conditions and high nitrogen concentrations, where the contribution from configurational entropy is low, binding energy effects will be predominant [46] ; hence they will promote the formation of the NV- complex, especially given its high binding energy ($E_{b}$ = 2.86 eV, calculated as $E_{b}=E_{f}(V_{Mg}^{2-})+E_{f}(N_{i}^{1+})-E_{f}(NV^{-})$; $V_{Mg}^{2-}$ and $N_{i}^{1+}$ denote the doubly-negatively charge magnesium vacancy and the singly-positively charged nitrogen interstitial). In general, the defects formed will strongly depend on the sample preparation techniques (for example, nitrogen incorporation during growth vs. implantation). Fig. 2a is indicative of the fact that the NV- defect can be formed in MgO by optimizing sample preparation conditions, as it is stable within the bandgap. Additionally, we checked for the possibility of the nitrogen interstitial recombining with the magnesium vacancy to form a substitutional nitrogen defect on the magnesium site. We found that the NV- defect is the preferred configuration; this result can be understood from a molecular orbital analysis, presented in Supplementary Note 3. Finally, we also computed the barrier for the nitrogen interstitial to migrate to other interstitial sites around the vacancy (see Supplementary Note 4). The computed barrier is high ($\approx 1.37$ eV), thus, once formed, the NV- defect is expected to be stable against annealing.
@@
Overall, the ground state properties of the NV- in MgO indicate that this defect is stable over a large energy range within the band gap of the host and thus, should be experimentally realizable. It has a spin triplet ground state with isolated occupied and unoccupied levels, pointing to the possibility of optical transitions to prepare and measure a qubit state. The computed ZFS tensor shows that a sufficiently large separation between spin sublevels is present, and that initialization in the $m_{s}$ = 0 sublevel should be realizable experimentally. Further, the computed hyperfine parameters indicate that the NV- center is expected to be distinguishable from other nitrogen defects in MgO.
@@ II.3 Excited state properties using hybrid TDDFT
We now turn to investigate the excited state properties of the NV- center using time-dependent density functional theory (TDDFT) with the DDH functional, as implemented in the WEST code [51]. We used spin-conserving and spin-flip TDDFT-DDH to investigate triplet and singlet excited states, respectively. The relaxed structure of the triplet ground state and first excited triplet state (obtained using DDH and TDDFT-DDH, respectively) are shown in Fig. 3a.
@@
Having determined the vertical excitation energies, we proceeded to relax the structure in the first excited triplet (${}^{3}E$) state using TDDFT-DDH with analytical forces [51] (see Fig. 3a). A decomposition of the many-body state shows that the $a_{1}$rightarrow $e$ excitations are dominant. We note that in the relaxed geometry of the excited state, the symmetry is reduced to $C_{1}$, with dN-Mg = 1.98 - 2.86 Åand dN-O = 1.37 Å. We find four equivalent minima, with the N-O bond tilted along the $+x$, $-x$, $+y$, $-y$ directions, consistent with the $C_{4v}$ symmetry of the ground state. Table 3 lists the absorption, zero phonon line (ZPL), and emission energies of the NV- in MgO (calculated under the Franck-Condon approximation) and compares it with that of the NV- center in diamond and the VV0 in silicon carbide. In MgO, the absorption and ZPL of the NV- center are in the ultra-violet (UV) range whereas the emission is in the visible range, highlighting a large Stokes shift.
@@
A large Stokes shift implies that a large number of phonons are involved in the emission process, which could lower the intensity of the ZPL relative to the rest of the emission spectrum, thus yielding an unfavorable Debye-Waller factor (DWF). The average number of phonons emitted during an electronic transition is called the Huang-Rhys factor (HRF, denoted as $S$) and is given by [58]
@@
We calculated $\omega_{eff}$, and therefore, the HRF $S$, by constructing a one-dimensional configuration coordinate diagram (CCD) along the direction connecting the relaxed ${}^{3}A_{2}$ ground and ${}^{3}E$ excited state structures [60, 57] (geometries optimized using DDH and TDDFT-DDH, respectively), as shown in Fig. 3c. By fitting a quadratic function around the points near the respective minima, we obtain the effective phonon frequencies of the ground and excited states to be $\hbar\omega_{g}$ = 32.79 meV and $\hbar\omega_{e}$ = 61.77 meV respectively. Using $\hbar\omega_{g}$ and $\Delta Q$, we obtain the HRF for emission as $S$ = 44.49, which is again indicative of a strong electron-phonon coupling [58], as expected. Such a large HRF leads to a negligible Debye-Waller factor (DWF $\simeq e^{-S}$approx 10-20), as a majority of the photons contribute to the emission side band instead of the ZPL. Thus, the ZPL will likely be undetectable, rendering the defect unsuitable as a single photon emitter (for example, for quantum communication applications). However, the defect’s triplet ground state, large ZFS, and many-body level structure may make it suitable for quantum sensing, computing, or transduction, by enabling optical initialization and spin-selective photo-luminescence. In fact, the possibility of optical initialization and readout is further highlighted from the CCD of the singlet excited states plotted in Fig. 3c. We see that at the ${}^{3}E$ minimum, ${}^{1}A_{1}$ is nearly degenerate with ${}^{3}E$, and ${}^{1}B_{1}$ and ${}^{1}B_{2}$ are lower in energy than the ${}^{3}E$ state. Thus, despite the large $\Delta Q$ in this system, the singlet states’ potential energy surfaces do not cross that of the triplet ${}^{3}E$ excited state (or are nearly degenerate with the ${}^{3}E$ minimum, as in the case for ${}^{1}A_{1}$), indicating that spin-selective photoluminescence via inter-system crossing may be possible. While the rates for these transitions would need to be calculated, this is an important prerequisite for the possibility of optical initialization and readout. Finally, nanostructuring or the additional presence of extended defects can be used to lower the HRF and therefore increase the intensity of the ZPL, as explored in SiC [61].
Using $\hbar\omega_{g}$, $\hbar\omega_{e}$, and $\Delta Q$, we show the emission spectrum of the NV- in Fig. 3d, which is nearly a Gaussian with a peak at the Franck-Condon energy of 1.43 eV, and a line-width of 420 meV. The emission spectrum thus lies in the visible red-infrared region, making it easy to detect. Moreover, a broad line-width is advantageous as it would allow multiple frequency filters to be used (for example, to filter out emissions from an underlying substrate) while retaining a significant fraction of the emission spectrum to be employed for detection. The radiative lifetime is 23 $\mu$s, calculated as the inverse of the radiative rate $\Gamma_{R}=\frac{n_{r}\mu_{\text{em}}^{2}(E_{\text{em}})^{3}}{3\pi\epsilon_{0}c^{3}\hbar^{4}}$, where $n_{r}$ is the refractive index of MgO (1.73), $\mu_{em}$ is the transition dipole moment for emission (0.035 $e$Å) calculated using TDDFT-DDH, and $E_{em}$ is the emission energy calculated using TDDFT-DDH (1.76 eV). This value is three orders of magnitude larger than those computed for the NV- in diamond [62] and VV0 in SiC [63], but is similar to those computed for spin defects in 2D transition metal dichalcogenides [64]. While it would be preferable to have a shorter radiative lifetime, the non-radiative lifetime is likely longer, as the ${}^{3}E$ and ${}^{3}A_{2}$ potential energy surfaces cross only at large barrier energies, and non-radiative rates are typically low for defects with transition energies larger than 1.5 eV [65]. Moreover, we note than the predicted coherence time $T_{2}$ for an electron spin defect in MgO is 600 $\mu$s [5], thus, the spin state could still be retained during optical initialization and readout. Nonetheless, cavity enhancement may be required due to other competing processes like inter-system crossing. Cavity enhancement of over 1000 has recently been reported for transitions in Er3+-doped MgO [66]. This result points to the encouraging possibility of cavity enhancement of the transitions in the NV- center in MgO as well.
@@
To summarize the computed excited state properties of the NV- in MgO, we find the presence of triplet excited and singlet shelving states, pointing at the possibility of an optical initialization cycle and spin-selective photo-luminescence. Further, our results show that the absorption and ZPL energies are in the UV range, whereas the emission energy is in the visible (red) range. The visible emission energy, together with the computed broad emission side-band, is a convenient feature for experimental detection in applications such as quantum sensing and transduction. We also find a large $\Delta Q$ and related Stokes shift, large HRF, and low DWF, indicating strong vibronic coupling. Identifying the reason behind this strong coupling is key to determining engineering strategies that can mitigate it, and to enhance our understanding of oxides as hosts for spin defects. We explore this in further detail below.